### v2

# Connected Speech-Based Cognitive Assessment in Chinese and English

Abstract
We present a novel benchmark dataset and prediction tasks for investigating approaches to assess cognitive function through analysis of connected speech. The dataset consists of speech samples and clinical information for speakers of Mandarin Chinese and English with different levels of cognitive impairment as well as individuals with normal cognition. These data have been carefully matched by age and sex by propensity score analysis to ensure balance and representativity in model training. The prediction tasks encompass mild cognitive impairment diagnosis and cognitive test score prediction. This framework was designed to encourage the development of approaches to speech-based cognitive assessment which generalise across languages. We illustrate it by presenting baseline prediction models that employ language-agnostic and comparable features for diagnosis and cognitive test score prediction. Unweighted average recall was 59.2% in diagnosis, and root mean squared error was 2.89 in score prediction.
keywords:
Speech biomarkers, neurodegenerative diseases, cognitive assessment, computational paralinguistics
1 Introduction
Cognitive problems such as memory loss, speech and language impairment, and reasoning difficulties occur frequently among older adults and often precede the onset of dementia syndromes. Due to the high prevalence of dementia and the costs this implies to health systems worldwide [1], research into cognitive impairment for the purposes of dementia prevention and early detection has become a priority in healthcare. There is a need for cost-effective and scalable methods for assessment of cognition and detection of impairment, from its most mild forms to severe manifestations of dementia. Speech is an easily collectable behavioural signal which reflects cognitive function and, therefore, could potentially serve as a digital biomarker of cognitive function, presenting a unique opportunity for application of speech technology [2].
We aim to assess speech as a behavioural marker of cognition in a global health context by investigating its application to the modelling of cognitive health indicators in two major languages, namely, Chinese and English. In this paper, we focus on prediction of cognitive test scores and diagnosis of mild cognitive impairment (MCI) in older speakers of Chinese and English, using samples of connected speech produced in picture description tasks. Our aim was to investigate approaches that are language independent or build on comparable features. To this end, we created, and are sharing with the research community, recorded speech from study participants doing picture description tasks along with clinical and neuropsychological test data.
This dataset has been used as a benchmark for speech processing and machine learning tasks that are relevant to the detection of cognitive decline through analysis of connected speech data. It formed the basis of the TAUKADIAL Challenge, at Interspeech 2024 (http://luzs.gitlab.io/taukadial/). We hope that this new resource will stimulate research on speech biomarkers in the speech, signal processing, machine learning and biomedical research communities, enabling them to test existing methods or develop novel approaches on a new, standardised dataset which will remain available to the community for future research and replication of results.
2 Background
The field of speech-based approaches to detecting cognitive decline has grown considerably over the last two decades, with a major focus on detecting dementia or Alzheimer’s dementia (AD) in comparison to a control (neurotypical or normal cognition, NC) group. A smaller proportion of studies has focused on MCI detection [2]. Most studies report accuracy figures without class balance, where accuracy is a biased measure. For example, [3] report a relatively high accuracy, 97.71%, in a highly imbalanced dataset while [4] and [5] report comparably lower accuracy, 62%, in a more balanced datasets. Similarly, using speech data generated from a cognitive assessment (picture description task), [6] obtained 85.4% accuracy using text-based features only on an imbalanced set of 268 participants. In contrast, [7] generated a subset of the same data (164 participants), balanced for class, gender, and age, and reported 78.7% accuracy with only acoustic features from standardised feature sets developed for computational paralinguistics.
Clinical tests such as the Mini-Mental State Examination (MMSE) are often part of these studies as mere data descriptors, rarely used in prediction. Some studies [8, 9, 10, 11] have used MMSE results as a baseline for classification, against which to compare the speech-based classifier, but very few available studies go beyond classification and use speech-based approaches to predict MMSE scores or other cognitive tests. However, there has been a shift of focus toward it in recent years. For instance, lexico-semantic features extracted from picture descriptions have been used in a model that was able to explain 51% of the variance of cognitive scores at the time of speech collection, and 56% of cognitive scores in a 12-month follow-up [12]. Other approaches have also been published, such as [13], which used BERT to predict MMSE scores from denoised speech recordings from picture description tasks and reported a root mean squared error (RMSE) of 3.76. Another study reported that acoustic features alone predict MMSE scores with a mean absolute error (MAE) of 5.66 and an $R^{2}$ of 0.125, with a linear regression analysis that improved by adding age, sex, and years of education to the model, yielding a MAE = 4.97 and $R^{2}$ = 0.261 [14] on the balanced dataset used by [7].
None of these studies addresses multilingual models where research is scarce and heterogeneous. A study on the AZTIAHO database reported accuracy scores between 60% and 93.79% using only ad hoc acoustic features. While this database contains samples in English, French, Spanish, Catalan, Basque, Chinese, Arabian, and Portuguese, it is small (40 participants) and remarkably age- and class-imbalanced [15].
Another multilingual study used English and Swedish speech samples generated through picture description tasks and word embeddings to train models that obtained classification accuracy of 63% for English and 72% for Swedish [16] in MCI diagnosis, and 75% ($F_{1}=0.77$) in AD AD/NC classification on 57 French and 550 English samples [17]. More recently, a signal processing grand challenge addressed the issue of generalising speech-based predictive models across two languages: Greek and English [18]. Differently from our experimental setting, theirs involved training of models in one language and testing on another. The top performing systems had classification accuracy between 69% to 87% (AD vs NC) and MMSE score prediction errors RMSE between 4.79 and 3.72. To the best of our knowledge, this study addresses a gap in the literature by combining multilingual speech analysis, MCI detection, and prediction of cognitive scores.
3 Data
Speech data are most often obtained from tasks embedded in neuropsychological batteries. Our dataset consists of Chinese and English speech samples collected while the speakers participated in picture description tasks conducted as part of cognitive assessments in clinical settings.
English-speaking participants were recruited from a community in the United States through print and online advertisements targeted to adults aged 60-90 with memory concerns. Eligible participants were at least 60 years old, spoke and understood English, had adequate hearing and vision to participate in a telehealth session, were stable on or not taking nootropic medications, and had a negative self-reported history of major psychiatric disorder or other medical disorder or illness that could cause cognitive decline (e.g., traumatic brain injury). Participants were classified as either NC or MCI. To be classified as MCI, a neuropsychologist determined that participants met the following National Institute on Aging-Alzheimer’s Association (NIA-AA) criteria  [19]: (a) self-reported a decline in cognition, (b) documented impairment in memory (produced a score greater than or equal to -1.5 SD on an objective measure), c) preserved functional independence (obtained a global score of less than or equal to 0.5 on the Clinical Dementia Rating Scale [20] - interview with a loved one), and (d) not demented. The University of Delaware Institutional Review Board approved data collection.
After providing informed consent, participants completed an assessment session via videoconferencing that lasted approximately 90 minutes. During this session, participants completed the discourse protocol and cognitive-linguistic battery with an assessor [21]. The discourse protocol tasks relevant to this project are: 1) the ”Cookie Theft” picture description task [22] elicited with the prompt, ”Please tell me everything you see going on in this picture”; 2) the ”Cat Rescue” picture [23] elicited with the prompt, ”Tell me a story with a beginning, a middle, and an end”; and 3) the Norman Rockwell print ”Coming and Going” [24] elicited with the same prompt as the Cat Rescue task. The cognitive-linguistic battery included the MoCA [25], whose scores were mapped to MMSE in this dataset following accepted practice [26]. The assessor used a standardised script and materials to deliver the discourse protocol and audio-recorded the administration using high-quality audio recording guidelines. The study data collection was managed using Research Electronic Data Capture [27] tools.
In the study used for collecting Chinese-language data, inclusion criteria were participants between 60 and 90 years old, with at least six years of education, and no history of neurological or psychiatric disorders. The neurologist evaluated participants with MCI according to the NIA-AA criteria. The evaluation was based on their CDR scores, which had a global score of 0.5, and brain magnetic resonance imaging (MRI) conducted within two years before recruitment, which showed atrophy in regions related to Alzheimer’s disease. Picture description tasks were employed to elicit connected speech, and responses were recorded the responses using a digital recorder. Participants described a set of three pictures depicting Taiwanese culture, with the instruction to report everything they observed in each one. The evaluators refrained from providing feedback but encouraged participants to elaborate if their responses were insufficient. Ethical approval was obtained from the Institutional Review Board of [REMOVED FOR DOUBLE-BLIND REVIEW] Hospital, and all participants signed a written informed consent document. color=yellow]SL: Anonymise this for submission
The full dataset (English and Chinese) was age- and gender-balanced to avoid bias in modelling. We ensured that the speech recordings met suitable audio quality standards for processing. Propensity score matching [28] was employed to generate an unbiased training set. The dataset was matched to scores defined in terms of the probability of an instance being treated as AD given covariates age and sex estimated through logistic regression, and matching instances were selected. All standardised mean differences for the covariates, standardised mean differences for squares, and two-way interactions between covariates were well below 0.1, indicating that the resulting set was adequately balanced.
The training set contained both Chinese and English samples with three picture descriptions per participant. The test set comprised recordings from different participants, with the same mix of languages and picture descriptions. Basic descriptive statistics of training and test set are shown in Table 1. Overall, there are 507 speech samples (261 Chinese and 246 English) with total duration of 528 minutes, ratio of training to test samples is just over 3:1. The dataset has been made available to the wider research community via DementiaBank [21].
Table 1: Dataset description, age, MMSE and duration expressed and mean years, score and seconds respectively. Numbers in brackets correspond to standard deviation and range.
MCI NC
Age 73.36 (6.14, 61-87) 71.85 (6.65, 61-87)
Men 39.2% ($n=87$) 38.2% ($n=63$)
Women 60.8% ($n=135$) 61.8% ($n=102$)
MMSE 25.84 (3.73, 13-30) 29.07 (1.08, 25-30)
Duration 58.92 (36.6, 12.7-240.9) 63.07 (33.9, 10.2-209.6)
Figure 1: General architecture for multilingual cognitive assessment based on recorded speech.
4 Cognitive assessment tasks
The benchmark presented in this paper encompasses the following tasks: (a) a classification task, where we aimed to create models to distinguish NC speech from MCI speech, and (b) a cognitive test score prediction (regression) task, where we created models to infer the subject’s MMSE scores based on connected (spontaneous) speech data.
The MCI classification task is evaluated through specificity ($\sigma$), sensitivity ($\rho$) and $F_{1}$ scores for the MCI category. These metrics are computed as follows: $\sigma=\frac{T_{N}}{T_{N}+F_{P}}$, $F_{1}=\frac{2\pi\rho}{\pi+\rho}$, where $\pi=\frac{T_{P}}{T_{P}+F_{P}}$, $\rho=\frac{T_{P}}{T_{P}+F_{N}}$, $N$ is the number of patients, $T_{P}$ is the number of true positives, $T_{N}$ is the number of true negatives, $F_{P}$ is the number of false positives and $F_{N}$ the number of false negatives. The balanced accuracy metric (unweighted average recall, UAR) is used for the overall ranking of this task’s results. It is defined as follows: $\operatorname{UAR}=\frac{\sigma+\rho}{2}$.
The MMSE regression task is assessed using the RMSE, defined as $\operatorname{RMSE}=\sqrt{\frac{\sum_{i=1}^{N}(\hat{y}_{i}-y_{i})^{2}}{N}}$, where $\hat{y}$ is the predicted MMSE score, $y$ is the patient’s actual MMSE score, and $\bar{y}$ is the mean score.
5 Modelling approach
As our goal is to explore models that generalise across languages, we aimed to create a single predictive model for each task which encompassed features extracted from both languages. Thus, the general architectures of our classification and regression systems are shown in Figure 1, where comparable features extracted from both languages are combined into a single predictive model.
5.1 Feature extraction
The feature extraction procedure aimed to identify speech features that could generalise well across the two languages. For acoustic features, we tested two different approaches: a traditional feature engineering approach with a feature set that has been found useful in emotion recognition and other computational paralinguistics tasks (eGeMAPs), and a self-supervised feature learning approach.
The eGeMAPs feature set comprises the F0 semitone, loudness, spectral flux, MFCC, jitter, shimmer, F1, F2, F3, alpha ratio, Hammarberg index, and slope V0 features, along with numerous statistical functions applied to these features. This results in a total of 88 features for every audio recording [29].
For self-supervised feature extraction we used the pre-trained model wav2vec, without fine tuning, and extracted features directly from raw audio [30].
To balance the duration of all audio recordings, we zero-padded the audio recordings for feature extraction. The features are extracted from the feature extractor layer. Then we applied a dropout layer, followed by a feature aggregation layer and another dropout layer. For dimensionality reduction, we used MaxPool1d layer (with a size of 42000, and a stride of 10,000). The result was used as input features for the multilayer perceptron (MLP) models. This resulted in 512 features per audio recording.
Finally, we extracted linguistic features that could be compared across languages. The recordings were first transcribed using automatic speech recognition (ASR) and part-of-speech tagged. Then the following features were calculated: number of tokens, number of types, type-to-token ratio, density (the ratio verbs, adjectives, adverbs, prepositions, and conjunctions to the total number of tokens), verb ratio, and pronoun ratio. To account for variability in pictures and descriptions, the number of tokens and number of types were z-score normalised. color=yellow]SL: We should also try adding comparable language features to these models. Perhaps this will improve the classification baselines. The regression baseline seems quite strong with wav2vec alone though.
5.2 Classification and regression
color=yellow]SL: We need to include (IS’24 checklist): “a) An explanation of evaluation metrics used, b) An explanation of how models are initialized, if applicable. c) Some measure of statistical significance of the reported gains or confidence intervals. d) A description of the computing infrastructure used and the average runtime for each model or algorithm (e.g. training, inference etc). e) The number of parameters in each model. Multi-layer of Perceptron (MLP) models were trained on different combinations of the above described feature sets using the Adam solver with relu activation. MLP models were employed for both classification and regression. We set $\alpha=10^{-4}$, hidden layers of sizes 55, 160, 160 and 55, constant learning rate 0f 0.001 and a maximum of 10,000 iterations. In both cases, 20-fold cross-validation was employed. The models were developed on an Intel Core i9-9980HK CPU @ 2.40GHz 2.40 GHz with 16 GB RAM and 8 GB GPU memory (Nvidia GeForce RTX 2080 with max-q design). The software used for balancing the dataset, feature extraction, model training, cross-validation and testing is available at https://gitlab.com/[REMOVEDFORDOUBLE-BLINDREVIEW]. color=yellow]SL: We need to add (as per Interspeech instructions: “A description of the computing infrastructure used and the average runtime for each model or algorithm (e.g. training, inference etc), and the number of parameters in each model. color=yellow]SL: Was hyperparameter search conducted. If so, include: f) Final results on a held-out evaluation set not used for hyperparameter tuning. g) Hyperparameter configurations for best-performing models. h) The method for choosing hyperparameter values to explore, and the criterion used to select among them. color=yellow]SL: Upload the source code to the gitlab repo and add a note here incidcating how the code can be obtained and the results reproduced.
6 Results
For the classification (diagnostic) task, our model achieved a test-data UAR of 59.18% while fusing the wav2vec and eGeMAPs features. The full set of results is shown in Table 2. Confidence interval were obtained through bootstrapping over 1000 runs [31]. The baseline result for this task is 59.18% UAR obtained on test data ($\sigma=0.587$, $\rho=0.597$, $\pi=0.617$). The results were very similar in both languages (English: UAR$=60.00\%$, $\sigma=0.40$, $\rho=0.80$; Chinese: UAR$=60.04\%$, $\sigma=0.39$, $\rho=0.81$). The overall accuracy score was 0.592, while F1 reached 0.602. Figure 2 shows the effect of each feature set on classification performance and Table 3 shows the results for both languages for the best performing methods. color=yellow]SL: The problem with these results is that if we were to choose a model purely based on CV we would have chosen the wav2vec, which clearly overfits, giving a test set UAR of just over 46%. We need to address this.
Table 2: Summary of results for the classification task (T.1), in % UAR, and the MMSE regression task (T.2), in RMSE for different features set combinations, where w2v = wav2vec and ling = comparable linguistic features, with confidence intervals in square brackets.
eGeMAPs w2v w2v+eGeMAPs linguistic w2v+linguistic ling+eGeMAPs hard fusion (all)
T.1 CV 66.17 [61.7, 71.4] 61.60 [56.7, 66.4] 50.94 [46.1, 55.7] 63.01 [59.6, 69.6] 59.08 [54.5, 64.1] 61.65 [56.8, 66.9] 66.22 [62.8, 72.5]
Test 54.89 [45.2, 63.4] 46.05 [33.3, 55.7] 59.18 [50.2, 68.7] 54.73 [46.1, 63.9] 51.71 [39.4, 65.1] 52.22 [42.6, 61.0] 53.26 [44.7, 63.2]
T.2 CV 4.02 [3.6, 4.5] 3.70 [3.3, 4.1] 3.82 [3.5, 4.2] 2.86 [2.5, 3.2] 3.44 [3.1, 3.8] 3.88 [3.5, 4.3] 3.04, [2.7, 3.4]
Test 3.82 [3.3, 4.3] 4.48 [4.1, 4.9] 3.76[3.2, 4.3] 2.89 [2.3, 3.5] 3.73 [3.3, 4.2] 3.45 [3.1, 3.8] 3.08, [2.7, 3.5]
Figure 2: Venn diagram showing the effect of each features set on classification with respect to Ground Truth (GT).
For the regression task, the comparable linguistic features on their own proved to be the most effective features, with RMSE scores of 2.86 (r = 0.514) and 2.89 (r = 0.337) for validation and test sets, respectively. Combining wav2vec and linguistic features also proved effective, but the eGeMAPs acoustic features were not found to be useful in this task. Unlike classification, regression results differed by language. For English, the RMSE was 1.75, while for Chinese the RMSE was 3.71, reflecting the standard deviations of MMSE (4.11 for Chinese and 1.27 for English).
7 Discussion
The present dataset is considerably less heterogeneous in terms of diagnoses and cognitive test scores than most public data used to date in research on predictive models for cognitive function assessment, including the few existing cross- and multi-lingual speech datasets used in this area [18, 16]. This makes the learning tasks defined in this paper harder, as they need to discriminate over a narrower range of values. However, our baseline models perform comparably to those models.
For the cognitive score prediction task (regression), we achieve an RMSE score of 2.89, while [13], for instance, reports an RMSE of 3.76, but using only the English subset of our data. The most comparable research is that conducted in a signal processing grand challenge to generalise speech-based predictive models across Greek and English [18, 32]. The best performing models achieved a classification accuracy between 69% and 87% (AD vs NC) and a RMSE between 4.79 and 3.72 for MMSE score prediction. However, these models involved training in one language and testing in another, while our experimental setup yields comparable results combining both languages in training and test, to our knowledge, for the first time.
Our goal was not to push the state-of-the-art on these datasets, but rather to establish proof-of-concept for our multilingual speech-only model’s capabilities to predict MMSE scores and detect MCI on a homogeneous multilingual dataset. Given that (a) our baseline results are comparable to other models in the literature (59.18%UAR and 2.89 RMSE), (b) both MMSE prediction and MCI detection are relatively uncommon compared to AD detection in the literature [2], (c) this is the first speech-only model built through combining datasets in two different languages, and (d) the models seem to generalise well for diagnosis across language (as suggested by the similarity in performance across the classification tasks), we argue that our work significantly contributes to the development of the field and will serve as a workable baseline for the wider research community.
As a limitation of this study, it should be noted that MMSE has been criticised for low discrimination (ceiling effect), especially in preclinical dementia [33]. This limitation is also common in similar studies in this research field. Therefore, future studies should aim to focus on other cognitive tests, more able to discriminate early stages of cognitive impairment.
A distinctive characteristic of our approach is the use of languages-agnostic and comparable languages-specific features. Our results suggest that comparable linguistic features can be valuable in MMSE prediction. While the fusion of wav2vec acoustic features to linguistic features did not improve on the results obtained with linguistic features alone, we believe that this approach should be explored further in larger datasets.inline,size=]ALL: here we will discuss our results, the limitations of the data and approaches, etc.
Fasih: can you add some comments about the feature sets, and in particular about the notably superior performance of wav2vec over the alternatives in MMSE prediction?
Table 3: Results Insights: Comparison of the best performing methods for the classification task (T.1) in % UAR and the MMSE regression task (T.2) in RMSE across both languages
T.1 T.2
English Cross-validation 52.7 1.43
Test 60.0 1.75
Chinese Cross-validation 48.3 3.73
Test 60.4 3.71
8 Conclusion
This paper presented a novel benchmark dataset for the development and testing of models for cognitive assessment through automatic analysis of connected speech. In particular, it defined learning tasks for diagnosis of MCI and prediction of MMSE scores. A general processing architecture for cross-lingual cognitive assessment was proposed which encompassed language-agnostic acoustic features and comparable linguistic features in a single predictive model for English and Chinese speech. Baseline models illustrated these predictive tasks and approach to feature extraction. The data and metadata have been made available to the research community. With the increasing interest by the medical community in speech biomarkers as a convenient and cost-effective approach to early detection and monitoring of cognitive problems, we expect this new resource will stimulate further research in the little explored field of cross-lingual modelling of cognitive function. color=yellow]SL: Please feel free add to and revise this section.
9 Acknowledgements
inline,size=]Acknowledgement should only be included in the camera-ready version, not in the version submitted for review. The 5th page is reserved exclusively for acknowledgements and references. No other content must appear on the 5th page.
SL acknowledges support from the British Academy/Leverhulme through Research Grant number SRG2223/231718.